# Graph-level Representation Learning with Joint-Embedding Predictive Architectures

### v2

@@ Abstract
Joint-Embedding Predictive Architectures (JEPAs) have recently emerged as a novel and powerful technique for self-supervised representation learning. They aim to learn an energy-based model by predicting the latent representation of a target signal $y$ from the latent representation of a context signal $x$. JEPAs bypass the need for negative and positive samples, traditionally required by contrastive learning while avoiding the overfitting issues associated with generative pretraining. In this paper, we show that graph-level representations can be effectively modeled using this paradigm by proposing a Graph Joint-Embedding Predictive Architecture (Graph-JEPA). In particular, we employ masked modeling and focus on predicting the latent representations of masked subgraphs starting from the latent representation of a context subgraph. To endow the representations with the implicit hierarchy that is often present in graph-level concepts, we devise an alternative prediction objective that consists of predicting the coordinates of the encoded subgraphs on the unit hyperbola in the 2D plane. Through multiple experimental evaluations, we show that Graph-JEPA can learn highly semantic and expressive representations, as shown by the downstream performance in graph classification, regression, and distinguishing non-isomorphic graphs. The code will be made available upon acceptance.
@@ 1 Introduction
Generative-based pretraining methods, on the other hand, typically remove or corrupt portions of the input and predict them using an autoencoding procedure (Vincent et al., 2010; He et al., 2022) or rely on autoregressive modeling (Brown et al., 2020; Hu et al., 2020). Fig. 1b depicts the typical instantiation of these methods: The input signal $x$ is fed into an encoder network that constructs the latent representation, and from it a decoder generates $\hat{y}$, the data corresponding to the target signal $y$. The energy function is then applied in data space, often through a reconstruction error. Generative models generally display strong overfitting tendencies van den Burg & Williams (2021) and can be non-trivial to train due to issues such as mode collapse Adiga et al. (2018). Moreover, the features they learn are not always useful for downstream tasks Meng et al. (2017). An intuitive explanation of this problem is that generative models have to estimate a data distribution that is usually quite complex, so the latent representations must be directly descriptive of the whole data space Loaiza-Ganem et al. (2022). This can become even more problematic for graphs because they live in a non-Euclidean and inhomogenous data space. Despite the aforementioned issues, masked autoencoding has recently shown promising results also in the graph domain with appropriately designed models (Hou et al., 2022; Tan et al., 2023).
@@
The nature of graph-level concepts is often assumed to be hierarchical (Ying et al., 2018). We conjecture that the typical latent reconstruction objective used in current JEPA formulations is not enough to provide optimal downstream performance. To this end, we design a prediction objective that starts by expressing the target subgraph encoding as a high-dimensional description of the hyperbolic angle. The predictor module is then tasked with predicting the location of the target in the 2D unit hyperbola. This prediction is compared with the target coordinates obtained by using the aforementioned hyperbolic angle. In this self-predictive setting, we explain why the stop-gradient operation and a simple predictor parametrization are useful to prevent representation collapse. To experimentally validate our approach, we evaluate Graph-JEPA against established contrastive and generative graph-level SSL methods across various graph datasets from different domains. Our proposed method demonstrates superior performance, outperforming most competitors while maintaining efficiency and ease of training. Notably, we observe from our experiments that Graph-JEPA can run up to 2.5x faster than Graph-MAE (Hou et al., 2022) and 8x faster than MVGRL (Hassani & Khasahmadi, 2020). Finally, we empirically demonstrate Graph-JEPA’s ability to learn highly expressive graph representations by showing that a linear classifier trained on the learned representations almost perfectly distinguishes pairs of non-isomorphic graphs that the 1-WL test cannot differentiate.
@@ 2.1 Self-Supervised Graph Representation Learning
Graph Neural Networks (GNNs)(Wu et al., 2019; Scarselli et al., 2008; Kipf & Welling, 2016a; Veličković et al., 2017) are now established solutions to different graph machine learning problems such as node classification, link prediction, and graph classification. Nevertheless, the cost of labeling graph data is quite high, given the immense variability of graph types and the information they can represent. To alleviate this problem, SSL on graphs has become a research frontier, where we can distinguish between two major groups of methods(Xie et al., 2022b; Liu et al., 2023):
@@
The goal of generative models is to recover the data distribution, an objective that is typically implemented through a reconstruction process. In the graph domain, most generative architectures that are also used for SSL are extensions of Auto-Encoder (AE) models (Hinton & Zemel, 1993) architectures. These models learn an embedding from the input data and then use a reconstruction objective with (optional) regularization to learn the data distribution. Kipf & Welling (2016b) extended the framework of AEs and VAEs (Kingma & Welling, 2013) to graphs by using a GNN as an encoder and the reconstruction of the adjacency matrix as the training objective. However, the results on downstream tasks with embeddings learned in this way are often unsatisfactory compared with contrastive learning methods, a tendency also observed in other domains (Liu et al., 2023). A recent and promising direction is masked autoencoding (MAE) (He et al., 2022), which has proved to be a very successful framework for the image and text domains. GraphMAE (Hou et al., 2022) is an instantiation of MAEs in the graph domain, where the node attributes are perturbed and then reconstructed, providing a paradigm shift from the structure learning objective of GAEs. S2GAE (Tan et al., 2023) is one of the latest GAEs, which focuses on reconstructing the topological structure but adds several auxiliary objectives and additional designs. Our architecture differs from generative models in that it learns to predict directly in the latent space, thereby bypassing the necessity of remembering and overfitting high-level details that help maximize the data evidence (Fig. 1b vs. 1c).
@@ 3 Method
A general overview.
We consider graphs $G$ defined as $G=(V,E)$ where $V=\{v_{1}\ldots v_{N}\}$ is the set of nodes, with a cardinality $|V|=N$, and $E=\{e_{1}\ldots e_{M}\}$ is the set of edges, with a cardinality $|E|=M$. For simplicity of the exposition, we consider symmetric, unweighted graphs, although our method can be generalized to weighted or directed graphs. In this setting, $G$ can be represented by an adjacency matrix $A\in\{0,1\}^{N\times N}$, with $A_{ij}=1$ if nodes $v_{i}$ and $v_{j}$ are connected and 00 otherwise. Fig. 2 provides the overview of the proposed architecture. The high-level idea of Graph-JEPA is to divide the input graph into subgraphs (patches) (He et al., 2023) and then predict the representation of a randomly chosen target subgraph from the representation of a single context subgraph (Assran et al., 2023). Again, we would like to stress that this masked modeling objective is realized in latent space without needing negative (or positive) samples. The subgraph representations are then pooled to create a vector representation for the whole graph, i.e., a graph-level representation. Therefore, the learning procedure consists of a sequence of operations: 1. Spatial Partitioning; 2. Subgraph Embedding; 3. Context and Target Encoding; 4. Latent Target Prediction.
Figure 2: A complete overview of Graph-JEPA. We first extract non-overlapping subgraphs (patches) (a.), perform a 1-hop neighborhood expansion (b.), and encode the subgraphs with a GNN (c.). After the subgraph encoding, one is randomly picked as the context and $m$ others as the targets (d.), and they are fed into their respective encoders (e.). The embeddings generated from the target encoder are used to produce the target subgraphs’ coordinates $\psi_{y}$. Finally, the predictor network is tasked with directly predicting the coordinates $\hat{\psi}_{y}$ for each target subgraph based on the context embedding and the positional embedding of each target subgraph (f.). A regression loss acts as the energy function $D$ between the predicted and target coordinates. Note that the extracted subgraphs in (a.) and (b.) are meant for illustrative purposes only. The number of nodes in each subgraph can vary.
@@ 3.1 Spatial Partitioning
We base the initial part of our architecture on the recent work of (He et al., 2023), but similar ideas consisting of graph partitioning have been proposed before for Graph SSL (Jin et al., 2020). This step consists of creating different subgraphs (patches), similar to how Vision Transformers (ViT) (Dosovitskiy et al., 2020) operate on images. We rely on the METIS (Karypis & Kumar, 1998) graph clustering algorithm, which partitions a graph into a pre-defined, non-overlapping number of clusters $p$ (Fig. 2a), such that the number of within-cluster links is much higher than between-cluster links. Note that having non-overlapping subgraphs can be problematic since edges can be lost in this procedure, and it is possible to end up with empty "subgraphs". In order to maintain a simple notion of locality in each subgraph and avoid completely empty subgraphs, a one-hop neighborhood expansion [^1] Connecting all (not already present) adjacent nodes from the input graph with each node present in the subgraph. of the nodes in each extracted subgraph is performed (Fig. 2b).
@@ 3.2 Subgraph Embedding
After partitioning the graph, we learn a representation for each subgraph through a GNN (Fig. 2c.). The specific choice of the GNN is arbitrary and depends on what properties one wishes to induce in the representation. The learned node embeddings are mean pooled to create a vector representation for each subgraph: $\{h_{1}...h_{p}\},h\in\mathbb{R}^{d}$. Given that these embeddings will be used as context or target variables, providing additional information regarding the subgraphs is key in order to guide the predictor network. Otherwise, the prediction task might be too difficult. Thus, we propose to use a positional embedding for each subgraph, which is implemented as the maximum Random Walk Structural Embedding (RWSE) of all the nodes in that subgraph. In this way, the position is characterized consistently for each patch. Formally, a RWSE (Dwivedi et al., 2021) for a node $v$ can be defined as:
$P_{v}=(M_{ii},M^{2}_{ii},\ldots,M^{k}_{ii})$ (1)
with $P_{v}\in\mathbb{R}^{k}$, $M^{k}=(D^{-1}A)^{k}$ the random-walk transition matrix of order $k$, and $i$ is the index of node $v$ in the adjacency matrix. Therefore, $M^{k}_{ii}$ encodes the probability of node $v$ landing to itself after a $k$-step random walk. Given a subgraph $l$, let $V_{l}$ denote the set of nodes in $l$. We define the subgraph RWSE as:
$P_{l}=\max\limits_{v\in V_{l}}P_{v}$ (2)
where the $\max$ operation is performed elementwise. A straightforward explanation of the above definition is that the positional information of each subgraph is essentially contained in the node with the highest degree [^2] A more suitable term would be the in-degree, but there is no difference in the undirected case., which will act as an anchor reference when predicting the target subgraphs’ latent representations. Intuitively, knowing $P_{l}$ is useful for the prediction task because the features present in the most representative node will have been diffused the most (via the GNN) into the subgraph.
@@ 3.3 Context and Target Encoding
Given the subgraph representations and their respective positional embeddings, we frame the Graph-JEPA prediction task in a similar manner to I-JEPA (Assran et al., 2023). The goal of the network is to predict the latent embeddings of randomly chosen target subgraphs, given one random context subgraph. The prediction is conditioned on positional information regarding each subgraph. At each training step, we choose one random subgraph as the context $x$ and $m$ others as targets $Y=\{y_{1},\ldots,y_{m}\}$ (Fig. 2d). These subgraphs are processed by the context and target encoders (Fig. 2e) which are parametrized by Transformer encoder blocks (without self-attention for the context) where normalization is applied at first (Xiong et al., 2020). The target encoder uses Hadamard self-attention (He et al., 2023), but other choices, such as the standard self-attention mechanism (Vaswani et al., 2017) are perfectly viable. We can summarize this step as:
@@
with $z^{x}\in\mathbb{R}^{d}$ and $Z^{y}\in\mathbb{R}^{m\times d}$. At this stage, we can use the predictor network to directly predict $Z^{y}$ from $z^{x}$. This is the typical formulation of JEPAs, also followed by the popular work of Assran et al. (2023). We argue that learning how to organize concepts for abstract objects such as graphs or networks directly in Euclidean space is suboptimal. In the following subsection, we propose a simple trick to bypass this problem using the encoding and prediction mechanisms in Graph-JEPA. A discussion in Section 4.3 will provide additional insights regarding this topic.

### v3

@@ Abstract
Joint-Embedding Predictive Architectures (JEPAs) have recently emerged as a novel and powerful technique for self-supervised representation learning. They aim to learn an energy-based model by predicting the latent representation of a target signal $y$ from the latent representation of a context signal $x$. JEPAs bypass the need for negative and positive samples, traditionally required by contrastive learning while avoiding the overfitting issues associated with generative pretraining. In this paper, we show that graph-level representations can be effectively modeled using this paradigm by proposing a Graph Joint-Embedding Predictive Architecture (Graph-JEPA). In particular, we employ masked modeling and focus on predicting the latent representations of masked subgraphs starting from the latent representation of a context subgraph. To endow the representations with the implicit hierarchy that is often present in graph-level concepts, we devise an alternative prediction objective that consists of predicting the coordinates of the encoded subgraphs on the unit hyperbola in the 2D plane. Through multiple experimental evaluations, we show that Graph-JEPA can learn highly semantic and expressive representations, as shown by the downstream performance in graph classification, regression, and distinguishing non-isomorphic graphs. The code is available at https://github.com/geriskenderi/graph-jepa.
@@ 1 Introduction
Generative-based pretraining methods, on the other hand, typically remove or corrupt portions of the input and predict them using an autoencoding procedure (Vincent et al., 2010; He et al., 2022) or rely on autoregressive modeling (Brown et al., 2020; Hu et al., 2020). Fig. 1b depicts the typical instantiation of these methods: The input signal $x$ is fed into an encoder network that constructs the latent representation, and from it a decoder generates $\hat{y}$, the data corresponding to the target signal $y$. The energy function is then applied in data space, often through a reconstruction error. Generative models generally display strong overfitting tendencies (van den Burg & Williams, 2021) and can be non-trivial to train due to issues such as mode collapse (Adiga et al., 2018). Moreover, the features they learn are not always useful for downstream tasks (Meng et al., 2017). An intuitive explanation of this problem is that generative models have to estimate a data distribution that is usually quite complex, so the latent representations must be directly descriptive of the whole data space (Loaiza-Ganem et al., 2022). This can become even more problematic for graphs because they live in a non-Euclidean and inhomogenous data space. Despite the aforementioned issues, masked autoencoding has recently shown promising results also in the graph domain with appropriately designed models (Hou et al., 2022; Tan et al., 2023).
@@
The nature of graph-level concepts is often assumed to be hierarchical (Ying et al., 2018). We conjecture that the typical latent reconstruction objective used in current JEPA formulations is not enough to provide optimal downstream performance. To this end, we design a prediction objective that starts by expressing the target subgraph encoding as a high-dimensional description of the hyperbolic angle. The predictor module is then tasked with predicting the location of the target in the 2D unit hyperbola. This prediction is compared with the target coordinates obtained using the aforementioned hyperbolic angle. In this self-predictive setting, we explain why the stop-gradient operation and a simple predictor parametrization are useful to prevent representation collapse. To experimentally validate our approach, we evaluate Graph-JEPA against established contrastive and generative graph-level SSL methods across various graph datasets from different domains. Our proposed method demonstrates superior performance, outperforming most competitors while maintaining efficiency and ease of training. Notably, we observe from our experiments that Graph-JEPA can run up to 2.5x faster than Graph-MAE (Hou et al., 2022) and 8x faster than MVGRL (Hassani & Khasahmadi, 2020). Finally, we empirically demonstrate Graph-JEPA’s ability to learn highly expressive graph representations by showing that a linear classifier trained on the learned representations almost perfectly distinguishes pairs of non-isomorphic graphs that the 1-WL test cannot differentiate.
@@ 2.1 Self-Supervised Graph Representation Learning
Graph Neural Networks (Wu et al., 2019; Scarselli et al., 2008; Kipf & Welling, 2016a; Veličković et al., 2017) are now established solutions to different graph machine learning problems such as node classification, link prediction, and graph classification. Nevertheless, the cost of labeling graph data is relatively high, given the immense variability of graph types and the information they can represent. To alleviate this problem, SSL on graphs has become a research frontier, where we can distinguish between two major groups of methods(Xie et al., 2022b; Liu et al., 2023):
@@
The goal of generative models is to recover the data distribution, an objective that is typically implemented through a reconstruction process. In the graph domain, most generative architectures that are also used for SSL are extensions of Auto-Encoder (AE) models (Hinton & Zemel, 1993) architectures. These models learn an embedding from the input data and then use a reconstruction objective with (optional) regularization to learn the data distribution. Kipf & Welling (2016b) extended the framework of different AEs and Variational AEs (Kingma & Welling, 2013) to graphs by using a GNN as an encoder and the reconstruction of the adjacency matrix as the training objective. However, the results on downstream tasks with embeddings learned in this way are often unsatisfactory compared with contrastive learning methods, a tendency also observed in other domains (Liu et al., 2023). A recent and promising direction is masked autoencoding (MAE) (He et al., 2022), which has proved to be a very successful framework for the image and text domains. GraphMAE (Hou et al., 2022) is an instantiation of MAEs in the graph domain, where the node attributes are perturbed and then reconstructed, providing a paradigm shift from the structure learning objective of GAEs. S2GAE (Tan et al., 2023) is one of the latest GAEs, which focuses on reconstructing the topological structure but adds several auxiliary objectives and additional designs. Our architecture differs from generative models in that it learns to predict directly in the latent space, thereby bypassing the necessity of remembering and overfitting high-level details that help maximize the data evidence (Fig. 1b vs. 1c).
@@ 3 Method
Notation and General Overview.
We consider graphs $G$ defined as $G=(V,E)$ where $V=\{v_{1}\ldots v_{N}\}$ is the set of nodes, with a cardinality $|V|=N$, and $E=\{e_{1}\ldots e_{M}\}$ is the set of edges, with a cardinality $|E|=M$. For simplicity of the exposition, we consider symmetric, unweighted graphs, although our method can be generalized to weighted or directed graphs. In this setting, $G$ can be represented by an adjacency matrix $A\in\{0,1\}^{N\times N}$, with $A_{ij}=1$ if nodes $v_{i}$ and $v_{j}$ are connected and 00 otherwise. Finally, let the neighborhood of a node $v$ be defined as $\mathcal{N}(v)=\{u\,|\,(u,v)\in E\}$. Fig. 2 provides the overview of the proposed architecture. The high-level idea of Graph-JEPA is to divide the input graph into subgraphs (patches) (He et al., 2023) and then predict the representation of a randomly chosen target subgraph from the representation of a single context subgraph (Assran et al., 2023). Again, we would like to stress that this masked modeling objective is realized in latent space without the need for negative samples. The subgraph representations are then pooled to create a vector representation for the whole graph, i.e., a graph-level representation. Therefore, the learning procedure consists of a sequence of operations: i) Spatial Partitioning; ii) Subgraph Embedding; iii) Context and Target Encoding; iv) Latent Target Prediction.
Figure 2: A complete overview of Graph-JEPA. We first extract non-overlapping subgraphs (patches) (a.), perform a 1-hop neighborhood expansion (b.), and encode the subgraphs with a GNN to learn feature vectors for the context and target (c.). Afterward, the context and target feature vectors are fed into their respective encoders (d.). The embeddings generated from the target encoder produce the target subgraphs hyperbolic coordinates $\psi_{y}$. On the other hand, the encoded context is fed into a predictor network, which is also conditioned on the positional embedding of the target subgraph, to then predict the coordinates $\hat{\psi}_{y}$ for the target subgraph (e.). A regression loss $D$, based on the distance in latent space, acts as the learning objective (f.). Note that the extracted subgraphs in (a.) and (b.) are meant for illustrative purposes only, as in practice, we use multiple target subgraphs for a given context. Furthermore, the number of nodes in each subgraph can vary.
@@ 3.1 Spatial Partitioning
We base the initial part of our architecture on the recent work of (He et al., 2023), but similar ideas have been proposed before for Graph SSL (Jin et al., 2020). This step consists of creating different subgraphs (patches), similar to how Vision Transformers (ViT) (Dosovitskiy et al., 2020) operate on images. Supported by experimental evidence in other data modalities (Assran et al., 2023; Fei et al., 2023), we conjecture that a critical aspect of training JEPAs is building a self-predictive task that is non-trivial and aligned with the characteristics of downstream problems. Therefore, we proceed by providing as input a graph partitioning, i.e., a collection of $p$ subsets, $\{V_{1},\ldots,V_{p}\}$ such that $V_{i}\cap V_{j}=\emptyset$ for $i\neq j$, $|V_{i}|=N/k$, and $\bigcup_{i}V_{i}=V$ (Fig. 2a). The rationale behind this choice is that the partition will inherently embed a notion of locality and spatial awareness into the learned representation via the latent self-predictive task, which is beneficial for downstream performance. In practice, we rely on the METIS (Karypis & Kumar, 1998) graph clustering algorithm, which partitions the graph such that the number of intra-cluster links is much higher than inter-cluster links, i.e., the number of edges whose incident vertices belong to different partitions is minimized. METIS is a multilevel algorithm, meaning that it first coarsens the graph, then performs partitioning, and finally, during an uncoarsening phase, projects the partitions back to the original graph, with their boundaries refined to reduce edge cuts and enhance intra-cluster density. This way, we can adequately embed local properties and community structure in the model input. It becomes pretty direct to see how these properties, especially the multilevel nature of the partitioning, relate to the hierarchical nature of graph-level concepts.
Note that having non-overlapping subgraphs can be problematic in practice since edges can be lost in this procedure, and it is possible to end up with edgeless subgraphs. Furthermore, signal propagation over the subgraphs becomes impossible due to the partitioning. Therefore, we implement a 1-hop neighborhood expansion (Fig. 2b), where each partition is transformed into $G_{i}^{\prime}=(V_{i}^{\prime},E_{i}^{\prime})$, such that $V_{i}^{\prime}=V_{i}\cap\mathcal{N}(V_{j}),\forall j\neq i$ and $E_{i}^{\prime}=\{(u,v)\,|\,u,v\in V_{i}^{\prime}\;\text{and}\;(u,v)\in E\}$. Simply put, all corresponding neighboring nodes from the original input graph are connected with each node in the subgraph. The model can then aggregate information from neighboring regions by incorporating these overlaps when learning subgraph embeddings. This enables it to capture broader graph-level patterns while retaining the properties of working with smaller, localized subgraphs. Such a property is crucial for the next step of Graph-JEPA, described in the following section.
@@ 3.2 Subgraph Embedding
After partitioning the graph, we learn a representation for each subgraph through a GNN (Fig. 2c.). The specific choice of the GNN is arbitrary and depends on what properties one wishes to induce in the representation. The learned node embeddings are mean pooled to create a vector representation for each subgraph: $\{h_{1}...h_{p}\},h\in\mathbb{R}^{d}$. These embeddings will be used as context or target variables in the JEPA framework (Fig. 1c.). Using only these embeddings can render the latent self-predictive too difficult, therefore, we include positional information for each subgraph. The positional embedding is implemented as the maximum Random Walk Structural Embedding (RWSE) of all the nodes in that subgraph. In this way, each patch’s position is characterized consistently and globally. Formally, a RWSE (Dwivedi et al., 2021) for a node $v$ can be defined as:
$P_{v}=(M_{ii},M^{2}_{ii},\ldots,M^{k}_{ii}),$ (1)
where $P_{v}\in\mathbb{R}^{k}$, $M^{k}=(D^{-1}A)^{k}$ is the random-walk transition matrix of order $k$, $D$ is a diagonal matrix containing the degrees of each node, and $i$ is the index $v$ in $A$. $M^{k}_{ii}$ encodes the probability of node $v$ landing to itself after a $k$-step random walk. It is noteworthy to see that such an embedding defines a signal over the graph, if we view the positional embedding as a function $\phi:V\rightarrow\mathbb{R}^{k}$. Furthermore, under a few reasonable assumptions, i.e., having connected, non-bipartite subgraphs with different connectomes, this function is injective $\phi(u)=\phi(v)\Rightarrow u=v,\;\forall u,v\in V$. This property follows directly from the definition of $M^{k}$ and the stationary distribution of a graph random walk Lovász (1993). Given a subgraph $l$, we define its RWSE as:
$P_{l}=\max\limits_{v\in V_{l}}P_{v},$ (2)
where the $\max$ operation is performed elementwise (on the nodes). From the above, we define the positional information of each subgraph as being essentially contained in the node with the highest degree [^1] The more suitable term would be in-degree, but there is no difference in the undirected case., which will act as an anchor reference when predicting the target subgraphs’ latent representations. Therefore, our choice of positional embedding will provide a global context for the predictor network, as the positional embeddings of these anchor nodes can identify the location of their respective subgraphs. Intuitively, knowing $P_{l}$ is useful for the prediction task because the features present in the most representative node will have been diffused the most (by the GNN) into the subgraph.
@@ 3.3 Context and Target Encoding
Given the subgraph representations and their respective positional embeddings, we frame the Graph-JEPA prediction task in a similar manner to I-JEPA (Assran et al., 2023). The goal of the network is to predict the latent embeddings of randomly chosen target subgraphs, given one random context subgraph, conditioned on the positional information of each target subgraph. At each training step, we choose one random subgraph as the context $x$ and $m$ others as targets $Y=\{y_{1},\ldots,y_{m}\}$. Using a single context is arbitrary but provides numerous advantages. Firstly, it guarantees that we can work with small and large graphs, as shown in Sec. 4.2. Secondly, it implicitly regularizes the model in learning long-range dependencies by predicting targets that are far (in terms of shortest path distance) from the context. We empirically validate this second aspect by testing for graph isomorphisms in Sec. 4.2. These subgraphs are processed by the respective context and target encoders (Fig. 2d), which are parametrized by Transformer encoder blocks (without self-attention for the context) where normalization is applied at first (Xiong et al., 2020). The target encoder uses Hadamard self-attention (He et al., 2023), but other choices, such as the standard self-attention mechanism (Vaswani et al., 2017), are perfectly viable. We can summarize this step as:
@@
with $z^{x}\in\mathbb{R}^{d}$ and $Z^{y}\in\mathbb{R}^{m\times d}$. At this stage, we can use the predictor network to directly predict $Z^{y}$ from $z^{x}$. This is the typical formulation of JEPAs, followed by the popular work of Assran et al. (2023). We argue that learning to organize concepts for abstract objects such as graphs or networks directly in Euclidean space is suboptimal. In the following subsection, we propose a simple trick to bypass this problem using the encoding and prediction mechanisms in Graph-JEPA. A discussion in Sec. 4.3 will provide additional insights.